OpenStudy (anonymous):

have an equilateral triangle with side length d. Two points given, A=(0,0) and B=(d,0). find point C of the triangle.

OpenStudy (anonymous):

 You know the middle point because it is between the other two, you also know the length of the base and the hypothenuse. Use that to discover the height.

OpenStudy (anonymous):

pythagorean theorem: (d/2)² + h² = d² (the hypothenuse) h² = d² - (d/2)² height \[h=\sqrt{d² - (\frac{ d }{ 2 })^{2}}\]

OpenStudy (anonymous):

And simply that (it comes out ... I want to say nice but I don't think you will think the same). Then you point is the middle over and the height up.

OpenStudy (anonymous):

\[h = \sqrt{d^{2}-\frac{ d^{2} }{ 4 }} = \sqrt{\frac{ 4d^{2} }{ 4 }-\frac{ 1d^{2} }{ 4 }}\] when taking power of fraction, take power on numerator and denominator?

OpenStudy (anonymous):

After simplifying that to \[\sqrt{\frac{ 3d ^{2} }{ 4 }}\] you can take the square root of the top and the bottom individually.

OpenStudy (anonymous):

\[h=\frac{ \sqrt{d^{2}3} }{ \sqrt{4} } = \frac{d\sqrt{3}}{2}\]and so \[C=(\frac{ d }{ 2 },\frac{d\sqrt{3}}{2})\]